Average Rate of Return—Cost Savings

Maui Fabricators Inc. is considering an investment in equipment that will replace direct labor. The equipment has a cost of $134,000 with a $12,000 residual value and a ten-year life. The equipment will replace one employee who has an average wage of $29,620 per year. In addition, the equipment will have operating and energy costs of $6,470 per year.

Determine the average rate of return on the equipment, giving effect to straight-line depreciation on the investment. If required, round to the nearest whole percent.
%
Answer #1

Depreciation = (134000-12000)/ 10= $12,200

Cost savings by purchase of equipment = Saving in labor cost - Depreciation - operating costs

= $29620-12200-6470

= $10,950

Average rate of return on investments = Cost savings by purchase of equipment / Average investment

= 10950 / (134000+12000)/2

= 15%